Verify that f(x)=^3square roor of 2x and g(x)=x^3/2 are inverse functions
Answer by RAY100(1637)

g(x) = x^(3/2)
let y=g(x)
y=x^(3/2)
inverse is x=y^(3/2)
raise both side to 2/3 power
x^(2/3) = y